OpenStudy (anonymous):

Find equation of curve which passes through (3,-4) and has slope 2y/x at the point (x,y) on it. I have the answer, just need to learn the logic...

OpenStudy (anonymous):

y+4=2y/x(x-3) y = 2y-6y/x -4 -y=-6y/x -4 y = 6y/x+4

OpenStudy (anonymous):

different from answer in book. for context, this is in a chapter called "Applications of first-order differential equations"

OpenStudy (anonymous):

the answer is 9y + 4x^2 = 0

OpenStudy (anonymous):

Anybody know how to get this answer?

OpenStudy (anonymous):

Looks a little peculiar, because if the curve is 9y + 4x^2 = 0, then dy/dx = -8x/9.

OpenStudy (anonymous):

Hmm, but -8x/9 does equal 2y/x, so I guess that's legit.

OpenStudy (anonymous):

Ok. yeah. Any advice on how to get to that answer?

OpenStudy (anonymous):

The way I would normally do it is to separate and integrate. If dy/dx = 2y/x, then 2dx/x = dy/y.

OpenStudy (anonymous):

that's the direction I went too. That got me to y = x^2 + c

OpenStudy (anonymous):

Ah, you're missing a constant multiple on y.

OpenStudy (anonymous):

Integrating 2dx/x = dy/y -> ln(x^2)=ln(y)+c exponentiate both sides ->x^2 = cy

OpenStudy (anonymous):

Can you get it from here @motza11 ?

OpenStudy (anonymous):

calculating...

OpenStudy (anonymous):

so integrating dy/y = 2/x dx, I get y = cx^2

OpenStudy (anonymous):

which is the same as yours since I can make "c" anything I want.. including 1/c...

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

now what, do I plug in (3,-4) to get the correct c?

OpenStudy (anonymous):

That's it. Then the last steps are to rearrange things to make it look pretty, and then to sit back and admire your work.

OpenStudy (anonymous):

BOOM!!!

OpenStudy (anonymous):

Thanks, Cliff!

OpenStudy (anonymous):

You're welcome. ;-)